OpenStudy (anonymous):

what are the base seven counting numbers just before and after 16 base seven ?

OpenStudy (anonymous):

just before is 15 just after is 20 In base 7 we have only the following digits: 0,1,2,3,4,5,6 where the 6 in base 7 is like the 9 in base 10.

OpenStudy (anonymous):

that's why in base 7, 20 comes after 16.

OpenStudy (anonymous):

THANKS FIDDLE

OpenStudy (anonymous):

SO 17 IS 7 AND 20 IS 14 ?

OpenStudy (anonymous):

I MEAN SO 10 IS 7 AND 20 IS 14

OpenStudy (anonymous):

yes, 10 (base 7) is 7 (base 10) 20 (base 7) is 14 (base 10)

OpenStudy (anonymous):

and 123 (base 7) = 1*7^2 + 2*7^1 + 3*7^0 = 49 + 14 + 3 = 66 (base 10)

OpenStudy (anonymous):

WHAT ARE THE LEAST AND GREATEST FOUR DIGIT NUMBERS AND THEIR DECIMAL EQUIVALENTS IN BASE 5?

OpenStudy (anonymous):

base 5 means we have the digits 0,1,2,3,4

OpenStudy (anonymous):

ONLY 0 1 2 3 4 R USED RIGHT?

OpenStudy (anonymous):

what would you say is the least one ?

OpenStudy (anonymous):

I WOULD SAY THE LEAST IS 0, BUT THAT IS NOT FOUR DIGITS.

OpenStudy (anonymous):

so, if you're not allowed to have leading zeros in the number - what would you pick as the smallest leading digit ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

and the entire number would be ?

OpenStudy (anonymous):

1000

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

now the largest 4 digit number ?

OpenStudy (anonymous):

IN BASE 5 WOULD BE 4444

OpenStudy (anonymous):

yep. now we need the decimal equivalents of the base 5 numbers 1000 and 4444 .

OpenStudy (anonymous):

???

OpenStudy (anonymous):

isn't that part of the question - to find the decimal equivalents of these two base 5 numbers ? do you know how to do it ?

OpenStudy (anonymous):

YES, BUT I DONT KNOW HOW?

OpenStudy (anonymous):

1000 (base 5) = 1 * 5^3 + 0 *5^2 + 0 *5^1 + 0 *5^0 = 1*125 + 0*25 + 0*5 + 0*1 = 125 (decimal)

OpenStudy (anonymous):

is what I did above clear ?

OpenStudy (anonymous):

I THINK SO, LET ME TRY TO FIGURE OUT THE GREATEST DECIMAL. GIVE ME A MIN

OpenStudy (anonymous):

624?

OpenStudy (anonymous):

can you post how you got that ?

OpenStudy (anonymous):

it's correct :)

OpenStudy (anonymous):

I DID WHAT U DID TO GET THE LEST DECIMAL. 4 * 5^3 + 4 * 5^2 + 4 *5^1 + 4 *5^0

OpenStudy (anonymous):

it's correct .

OpenStudy (anonymous):

WOULD I USE THE SAME FORMULA TO FIND THE DECIMAL EQUIVALNT FOR 65 BASE 9

OpenStudy (anonymous):

It's the same idea, just use 9^.. instead of 5^..

OpenStudy (anonymous):

9 * 6^2 + 9 * 5^1 ?

OpenStudy (anonymous):

6*9^1 + 5*9^0

OpenStudy (anonymous):

BUT 6 IS IN THE TENS PLACE, AND 5 IS IN THE ONES PLACE. WHY USE THE ZERO POWER IF IN THE ONES PLACE?

OpenStudy (anonymous):

9^0 = 1 so this represents the ones place. 9^1 = 9 which represents the "tens" place in base 9

OpenStudy (anonymous):

just like in base 10 (decimal) 10^0 = 1 represents the one's place 10^1 = 10 represents the tens place

OpenStudy (anonymous):

SO 65 BASE NINE IS DECIMAL # 59?

OpenStudy (anonymous):

yep that's right

OpenStudy (anonymous):

IF I USE ^1 FOR THE TENS AND 0^ FOR THE ONES......WHAT DO U USE FOR 0?

OpenStudy (anonymous):

to represent the number 0 ?

OpenStudy (anonymous):

THE ZERO PLACE IN A NUMBER.....LIKE 400?

OpenStudy (anonymous):

you are referring the the values of the digits being zero , not their place. In 400 , one zero is placed in the "ones" place and the other is placed in the "tens" place.

OpenStudy (anonymous):

OIC

OpenStudy (anonymous):

don't forget to click "good answer" :)

OpenStudy (anonymous):

SORRY

OpenStudy (anonymous):

np

OpenStudy (anonymous):

SO HOW WOULD IT WORK THE OTHER WAY AROUND, HOW DO I CONVERT 98 TO BASE FIVE?

OpenStudy (anonymous):

First you would write out the powers of the base you want to convert to (five). 5^0=1 5^1=5 5^2=25 5^3=125 And I'm stopping here, coz 125 is bigger than 98. Now, I start with 5^2 , and check how many times it goes into 98. 98 / 25 = 3 with remainder of 23 ( coz 98 - 3*25 = 23) This means, my answer will have 3 in the base 5 hundreds place (5^2). Next, I go to 5^1 , and check how many times it goes into 23 (remainder from above) 23/5 = 4 with remainder of 3 (coz 23 - 4*5 = 3) This means, my answer will have 4 in the base 5 tens place (5^1) Now, I am at 5^0 (the base 5 ones place). It will contain 3 (the remainder from the previous step). So, 98 (decimal) is 343 (base 5) To check if I got it right, I'll convert 343 ( base 5) back to decimal: 3*5^2 + 4*5^1 + 3 = 3*25 + 4*5 + 3 = 75 + 20 + 3 = 98 (decimal)

OpenStudy (anonymous):

it's a bit more annoying :)

OpenStudy (anonymous):

4479 TO BASE SIXTEEN. ?

OpenStudy (anonymous):

Start by listing base 16 powers: 16^0=0 16^1=16 16^2=256 16^3=4096 16^4=65536 <--- larger than 4479 so stopping here. The base 16 number will have 4 digits. We start with 16^3 . How many times does it go into 4479 , and what is the remainder ?

OpenStudy (anonymous):

1 R =383

OpenStudy (anonymous):

yes. that means that in the base 16 thousands position we will have a 1. Now we move to the base 16 hundreds place , with that remainder of 383 , and do the same thing.

OpenStudy (anonymous):

how many times does 16^2 go into 383 , and what is the remainder ?

OpenStudy (anonymous):

1 R 127

OpenStudy (anonymous):

so, we will have a 1 in the base 16 hundreds place. now take the remainder of 127 - and proceed with it to the base 16 tens place.

OpenStudy (anonymous):

how many times does 16^1 go into 127, and what is the remainder ?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

and remainder ?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

yes. so the 7 will go in the base 16 tens place . And the remainder 15 will go to the ones place.

OpenStudy (anonymous):

In base 16 we have the following digits: 0,1,2,3,4,5,6,7,8,9,A,B,C,D,E,F So our remainder of 15 (decimal) is represented by the digit F in base 16.

OpenStudy (anonymous):

117F

OpenStudy (anonymous):

So what would be our entire base 16 number ?

OpenStudy (anonymous):

yep. that's it.

OpenStudy (anonymous):

you can try converting it back to decimal to check it.

OpenStudy (anonymous):

A BOOKSTORE ORDERED 2 GROSS, 11 DOZEN, AND 7 PENCILS. EXPRESS IN BASE 12 AND BASE 10?????????